OpenStudy (wade123):

@ganeshie8 please help!

OpenStudy (wade123):

@freckles can you please pleaese explain??

OpenStudy (freckles):

\[\int\limits_{a}^{b}f(x) \approx (b-a) \cdot \frac{f(a)+f(b)}{2}\] I think that is the trapezoid approximation you are refering to

OpenStudy (freckles):

you are given a=1 and b=2 f(a)=f(1)=1 f(b)=f(2)=14

OpenStudy (freckles):

or you might be talking about this one: \[\frac{b-a}{2 N}(fx_1)+2f(x_2)+2f(x_3)+ \cdots+2f(x_N)+f(x_{N+1}))\]

OpenStudy (freckles):

So your N+1=7 since that is how many function values you are given So N=7-1=6

OpenStudy (freckles):

do you have any questions I think you are looking for the second one there I mentioned

OpenStudy (freckles):

x_1=1 x_2=1.1 x_3=1.3 x_4=1.6 x_5=1.7 x_6=1.8 x_7=2

OpenStudy (wade123):

ahhh wait can you do the steps with me? i dont realy understand the trapezoidapproximation

OpenStudy (freckles):

it is just a pluggin in thing doing using order of operations

OpenStudy (freckles):

\[\frac{b-a}{2 N}(f(x_1)+2f(x_2)+2f(x_3)+ \cdots+2f(x_N)+f(x_{N+1})) \\ \frac{2-1}{2(7-1)}(f(1)+2f(1.1)+2f(1.3)+2f(1.6)+2f(1.7)+2f(1.8)+f(2))\]

OpenStudy (wade123):

should i do the first or second one?

OpenStudy (freckles):

plug in and follow the order of operations

OpenStudy (freckles):

i assume the second one but you tell me

OpenStudy (freckles):

which one is called the trapezoid approximation in your class

OpenStudy (wade123):

do they both come out with the same answer? ill look in my book

OpenStudy (freckles):

Not always I bet.

OpenStudy (wade123):

okay one second, ill check

OpenStudy (freckles):

actually that rule looks a little different from the second one

OpenStudy (freckles):

that one has no 2's in front of the x_i's between x_1 and x_(N+1)

OpenStudy (freckles):

i mean in front of the f(x_i)'s between f(x_1) and f(x_(N+1)

OpenStudy (freckles):

\[T_n=\frac{\Delta x }{2}(y_0+y_1+y_2+ \cdots+y_n)\] So add up all the y values according to the rule you gave me and then multiply that sum by (b-a)/n that is (2-1)/7

OpenStudy (freckles):

oops (2-1)/(7(2))

OpenStudy (freckles):

forgot about the 2 on bottom (you know below the delta x)

OpenStudy (freckles):

\[\frac{b-a}{2 N}(f(x_1)+2f(x_2)+2f(x_3)+ \cdots+2f(x_N)+f(x_{N+1})) \\ \frac{2-1}{2(7-1)}(f(1)+2f(1.1)+2f(1.3)+2f(1.6)+2f(1.7)+2f(1.8)+f(2)) \\ \frac{1}{2(6)}(1+2(3)+2(5)+2(8)+2(10)+2(11)+14) \\ \frac{1}{12}(1+6+7+16+20+22+14) \\ \frac{1}{12}(7+23+42+14) \\ \frac{1}{12}(30+56) \\ \frac{1}{12}(86) \\ \frac{86}{12}=\frac{43}{6} \approx 7.2\] The other rule you provided should probably get something kinda close to this number

OpenStudy (wade123):

okay, thank you !!

OpenStudy (freckles):

so i guess we should have did 86/(7*2) here 86/14 if we want to use the same rule throughout your work